Find the surface-area-to-volume ratio for a rectangular prism with edge lengths 9,6,and 7.
Answer by rajagopalan(148)

The surface-area 0f prism 9,6,and 7
= 2(LB+BH+LH)=2(9*6 + 6*7 +7*9)
=2(54+42+63)=2(159)=318 sq units
volume for a rectangular prism with edges 9,6,and 7=9*7*6=63*6=378 Cu units
Ratio of Surface area : volume=318:378
=53:63
Ans 53:63
okay